A copper 20cm on a side is completely submerged underwater. Calculate the buoyant force acting on the cube?

a.58.7 N b. 75.8 N c. 78.5 N d. 87.5 N




1
Expert's answer
2021-04-26T17:00:18-0400

The volume of the cube is "V = l^3 = (0.2 m)^3 = 0.008 m^3".

Buoyant force, acting on it is "F = \rho g V", where "\rho" is the density of water, "g" is gravitational acceleration and "V" is the submerged volume.

Hence, "F = 1000 \frac{kg}{m^3} \cdot 9.81 \frac{m}{s^2} \cdot 0.008 m^3 \approx 78.5 N", so the answer is c).